# Getting it Right: Improving Spatial Consistency in Text-to-Image Models

### v1

@@ 5.3 Investigating the CLIP Text Encoder
The CLIP Text Encoder enables semantic understanding of the input text prompts in the Stable Diffusion model. As we fine-tune CLIP on the spatial captions from SPRIGHT, we investigate the various nuances associated with it:
Figure 8: Comparing layer-wise learned representations between the Baseline CLIP (X-axis) and its fine-tuned version on SPRIGHT (Y-axis). When compared on spatial captions, we find that distinct representations are learned in the output attention projections and the MLP layers, while layer norm layers higher levels of similarities. The gap in the learned representations increases when complex prompts are used, indicating that longer spatial prompts in SPRIGHT indeed lead to diverse representations in the CLIP embedding space.
Centered Kernel Alignment: The CKA metric [43, 30] allows us to compare the different layer-wise representations learned by two neural networks. Figure 8 illustrates different representations learned by baseline CLIP, compared against the one trained on SPRIGHT. We compare layer activations across 50 simple and complex prompts and aggregate representations from all the layers. Our findings reveal that the MLP and output attention projection layers play a larger role in enhancing spatial comprehension, as opposed to layers such as the layer norm. This distinction becomes more pronounced with complex prompts, emphasizing that the longer prompts from SPRIGHT lead to more diverse embeddings being learned within the CLIP space.
[…]
above 0.9225 0.8674
below 0.9259 0.8673
to the left of 0.9229 0.8658
to the right of 0.9223 0.8528
in front of 0.9231 0.8417
behind 0.9289 0.8713
Table 11: We find that CLIP fine-tuned on SPRIGHT is able to differentiate the spatial nuances present in a textual prompt. While Baseline CLIP shows a high similarity for spatially different prompts, SPRIGHT enables better fine-grained understanding.
@@
: To evaluate semantic understanding of the fine-tuned CLIP, we conduct the following experiment: Given an original prompt containing a spatial phrase and 2 objects, we modify the prompt by switching the objects. For example, an original prompt - "an airplane above an apple" is modified to "an apple above an airplane". Our rationale is that, although both sentences are semantically close to each other, they are spatially distinct. we posit that a better model should be able to discern the fine-grained spatial distinction that are conveyed through both the sentences. We perform this experiment for 6 kinds of spatial relationships, for a total of $\sim$ 37k sentences. For each model, we compute the cosine similarity between the pooled layer outputs of the original and modified prompts. We find, as shown in Table 11, that fine-tuned CLIP indeed is able to differentiate better between the spatial prompts. And, it does not suffer from catastrophic forgetting, which would have yielded a very low cosine similarity.
@@ 5.4 Training with Negation
5.5 Improvement over Attention Maps
Finally, inspired from methods such as Attend-and-Excite [5], we visualize attention relevancy maps for both simple and complex spatial prompts and find that we are able to better generate the expected object as well as achieve better spatial localization. In both examples, there are missing objects in the baseline while our model is able to generate them; the bed and the house are not present for Stable Diffusion 2.1 while our model succeeds in generating them. The relevancy map shows that the patches of the image that have high attention to the missing words is spread across the image. We can also see from this figure that the spatial word has high attention to the correct part of the image in our model, compared to the baseline: In our model, as shown in Figure 9 (bottom row), below attends to image patches that are below the bed and right attends to the image patches that are located on the road and on its right, while the Stable diffusion 2.1 model fails to give high relevancy map to the expected locations. We are able to improve upon both the aspects through the intermediate attention maps and the final generated images.
@@ Acknowledgements
We thank Lucain Pouget for helping us in uploading the dataset to the Hugging Face Hub and the Hugging Face team for providing computing resources to host our demo. The authors acknowledge resources and support from the Research Computing facilities at Arizona State University. This work was supported by NSF RI grants #1750082 and #2132724. The views and opinions of the authors expressed herein do not necessarily state or reflect those of the funding agencies and employers.
Appendix
Appendix A System Prompt for GPT-4 Evaluation
You are part of a team of bots that evaluates images and their captions. Your job is to come up with a rating between 1 to 10 to evaluate the provided caption for the provided image. Consider the correctness of spatial relationships captured in the provided image. Return the response formatted as a dictionary with two keys: ‘rating’, denoting the numeric rating, and ‘explanation’, denoting a brief justification for the rating.
The captions you are judging are designed to stress-test image captioning programs, and may include:
1. Spatial phrases like above, below, left, right, front, behind, background, foreground (focus most on the correctness of these words).
2. Relative sizes between objects such as small & large, big & tiny (focus on the correctness of these words).
3. Scrambled or misspelled words (the image generator should produce an image associated with the probable meaning). Make a decision as to whether or not the caption is correct, given the image.
A few rules:
1. It is ok if the caption does not explicitly mention each object in the image; as long as the caption is correct in its entirety, it is fine.
2. It is also ok if some captions don’t have spatial relationships; judge them based on their correctness. A caption not containing spatial relationships should not be penalized.
3. You will think out loud about your eventual conclusion. Don’t include your reasoning in the final output.
4. Return the response formatted as a Python-formatted dictionary having two keys: ‘rating’, denoting the numeric rating, and ‘explanation’, denoting a brief justification for the rating.
Appendix B Results on T2I-CompBench
Method Attribute Binding Object Relationship
Color Shape Texture Spatial Non-Spatial
SD 1.4 0.3765 0.3576 0.4156 0.1246 0.3079
SD 2 0.5065 0.4221 0.4922 0.1342 0.3096
Composable v2 [39] 0.4063 0.3299 0.3645 0.0800 0.2980
Structured v2 [14] 0.4990 0.4218 0.4900 0.1386 0.3111
Attn-Exct v2 [5] 0.6400 0.4517 0.5963 0.1455 0.3109
GORS [23] 0.6603 0.4785 0.6287 0.1815 0.3193
DALLE-2 [53] 0.5750 0.5464 0.6374 0.1283 0.3043
SDXL [49] 0.6369 0.5408 0.5637 0.2032 0.3110
PixArt-Alpha [6] 0.6886 0.5582 0.7044 0.2082 0.3179
Kandisnky v2.2 [26] 0.5768 0.4999 0.5760 0.1912 0.3132
DALL-E 3 [45] 0.8110 0.6750 0.8070 - -
Ours (<500 images) 0.6251 0.4648 0.5920 0.2133 0.3132
Table 12: Results on the T2ICompBench Benchmark. a) We achieve state of the art spatial score, across all methods, by efficient fine-tuning on only 444 images. b) Despite not explicitly optimizing for them, we find substantial improvement and competitive performance on attribute binding and non-spatial aspects.
As shown in Table 12, we achieve state of the art performance on the spatial score in the widely accepted T2I-CompBench [23] benchmark. The significance of training on images containing a large number of objects is emphasized by the enhanced performance of our models across various dimensions in T2I-CompBench. Specifically, we enhance attribute binding parameters such as color and texture, alongside maintaining competitive performance in non-spatial aspects.
Appendix C Illustrative Examples from VISOR
In Figure 10, we present illustrative results on the VISOR benchmark.
Appendix D Comparing COCO-30K and Generated Images
In Figure 11, we compare images from COCO, baseline Stable Diffusion and our model. We find that the generated images from our model adhere to the input prompts better, are more photo-realistic in comparison to the baseline.
Appendix E Additional Examples from SPRIGHT
Figure 12 and 13 demonstrate a few correct and incorrect examples present in SPRIGHT. In Figure 13, while most relationships are accurately described in the captions, there are instances where the model struggles to capture the precise spatial nuance.
Appendix F Contributions
AC defined the scope of the project, performed the initial hypothesis experiments and conducted the evaluations. GBMS led all the experimental work and customized the training code. EA generated the dataset, performed the dataset and relevancy map analyses. SP took part in the initial experiments, suggested the idea of re-captioning and performed few of the evaluations and analyses. DG suggested the idea of training with object thresholds and conducted the FAITHScore and GenEval evaluations. TG participated in the early discussions on spatial failures of T2I models. VL, CB, and YZ co-advised the project, initiated and facilitated discussions, and helped shape the the goal of the project. AC and SP wrote the manuscript in consultation with TG, LW, HH, VL, CB, and YZ. All authors discussed the result and provided feedback for the manuscript.
Appendix G Limitations
Since SPRIGHT is a derived dataset, it inherits the limitations of the original datasets. We refer the readers to the respective papers that introduced the original datasets for more details. As shown in our analysis, the generated synthetic captions are not a 100% accurate and could be improved. The improvements can be achieved through better prompting techniques, larger models or by developing methods that better capture low-level image-text grounding. However, the purpose of our work is not to develop the perfect dataset, it is to show the impact of creating such a dataset and its downstream impact in improving vision-language tasks. Since our models are a fine-tuned version of Stable Diffusion, they may also inherit their limitations in terms of biases, inability to generate text in images, errors in generating correct shadow patterns. We present our image fidelity metrics reporting FID on COCO-30K. COCO-30K is not the best dataset to compare against our images, since the average image resolutions in COCO is lesser than those generated by our model which are of dimension 768. Similarly, FID largely varies on image dimensions and has poor sample complexity; hence we also report numbers on the CMMD metric.
Figure 10: Generated images from our model, as described in Section 4.2, on evaluation prompts from the VISOR benchmark.
Figure 11: Illustrative examples comparing ground-truth images from COCO and generated images from Baseline SD 2.1 and our model. The images generated by our model exhibit greater fidelity to the input prompts, while also achieving a higher level of photorealism.
Figure 12: Illustrative examples from the SPRIGHT dataset, where the captions are correct in its entirety; both in capturing the spatial relationships and overall description of the image. The images are taken from CC-12M and Segment Anything.
Figure 13: Illustrative examples from the SPRIGHT dataset, where the captions are not completely correct. The images are taken from CC-12M and Segment Anything.

### v2

@@ 5.3 Investigating the CLIP Text Encoder
The CLIP Text Encoder enables semantic understanding of the input text prompts in the Stable Diffusion model. As we fine-tune CLIP on the spatial captions, we investigate the various nuances associated with it:
Centered Kernel Alignment
(CKA) [38, 28] compares layer-wise representations learned by two neural networks. Figure 5 illustrates different representations learned by baseline CLIP, compared against the one trained on SPRIGHT. We compare layer activations across 50 simple and complex prompts and aggregate representations from all the layers. Our findings reveal that the MLP and output attention projection layers play a larger role in enhancing spatial comprehension, as opposed to layers such as the layer norm. This distinction is larger with complex prompts, showing that the longer prompts from SPRIGHT indeed lead to more diverse embeddings being learned within the CLIP space.
@@
: To evaluate semantic understanding of the fine-tuned CLIP, we perform the following experiment: given a prompt containing a spatial phrase and 2 objects, we modify the prompt by switching the objects (e.g. “an airplane above an apple” $\rightarrow$ “an apple above an airplane”). Although these sentences have the same words, the placement of the two nouns relative to the preposition “above” completely changes the meaning of the sentence. To evaluate if models can discern this spatial distinction, we compute the cosine similarity between the pooled layer outputs of the original and modified prompts, for $\sim$ 37k sentences. Table 9 shows that CLIP finetuned on SPRIGHT is able to differentiate between the prompts better (i.e. lower cosine similarity) than the baseline.
5.4 Improvement over Attention Maps
Inspired by methods like Attend-and-Excite [5], we visualize attention relevancy maps for both simple and complex spatial prompts. Our model better generates the expected objects and achieves improved spatial localization compared to the baseline. For instance, the baseline models fails to generate objects like the bed and house, which our model successfully generates. The relevancy map indicates that high attention patches for missing words are spread across the image. Additionally, our model correctly attends to spatial words in the image, unlike the baseline. For example, in our model (Figure 6, bottom row), below attends to patches below the bed, and right attends to patches on the road’s right, while Stable Diffusion 2.1 does not. We achieve these improvements across the intermediate attention maps and the final generated images.
@@ Acknowledgements
We thank Lucain Pouget for helping us in uploading the dataset to the Hugging Face Hub and the Hugging Face team for providing computing resources to host our demo. The authors acknowledge resources and support from the Research Computing facilities at Arizona State University. AC, CB, YY were supported by NSF Robust Intelligence program grants #1750082 and #2132724. TG was supported by Microsoft’s Accelerating Foundation Model Research (AFMR) program and UMBC’s Strategic Award for Research Transitions (START). The views and opinions of the authors expressed herein do not necessarily state or reflect those of the funding agencies and employers.